A piece of alloy has a measured mass of 86g in the air and 73g when immersed in water. Find it's volume and it's density


1
Expert's answer
2021-08-31T08:58:20-0400

The weight of the alloy in the air is:


"P_{air} = mg = \rho Vg"

where "m = 86g = 0.086kg" is its real mass, "\rho" is its density and "V" is its volume. "g\approx 9.81m/s^2" is the gravitational acceleration.

In the water due to the buoyancy the weight is:


"P_{water} = P_{air} - F_b = m_{water}g"

where "m_{water} = 73g = 0.073kg" is the "effective" mass measured in water.

The buoyancy force is given as follows:


"F_b = \rho_{water}Vg"

where "\rho_{water} = 1000kg/m^3" is the density of water.

Thus, obtain the volume:


"mg - \rho_{water}Vg = m_{water}g\\
m - \rho_{water}V = m_{water}\\
V = \dfrac{m - m_{water}}{\rho_{water}}\\
V = \dfrac{0.086kg - 0.073kg}{1000kg/m^3} = 1.3\times 10^{-5}m^3"

And the density:


"\rho = \dfrac{m}{V} = \dfrac{0.086kg}{1.3\times 10^{-5}m^3} \approx 6615\space kg/m^3"



Answer. Volume: "1.3\times 10^{-5}m^3", density: "6615\space kg/m^3".